# Compositional Difference-in-Differences for categorical outcomes

### v2

@@ 1.3. Geometric justifications of parallel growths
In this section, I provide some geometric justifications for parallel growths. I start by interpreting the restriction on the shares using the compositional difference operator.
@@
Geometric interpretation Parallel growths also imply parallel trajectories of probability mass functions in the probability simplex. Probability mass functions lie in the simplex $\mathcal{S}^{p-1}$, where movement can be meaningfully defined using the geometry introduced by Aitchison (1982). Unlike the Euclidean case, this geometry accounts for the simplex’s constraints and forms the basis of compositional data analysis, redefining operations such as addition, scaling, and distance.
@@
. Let $\mathcal{S}^{p-1}$ be the p1p-1 simplex, define the following operations: for all $\pi_{1},\pi_{2}\in\mathcal{S}^{p-1}$, we have:
@@
Figure 3. Illustration of parallel growths in the simplex. The red curve represents the trajectory of the control group from pre-treatment ($\pi_{0,0}^{N}$) to post-treatment ($\pi_{0,1}^{N}$), while the blue curve shows the counterfactual trajectory of the treated group from pre-treatment ($\pi_{1,0}^{N}$) to post-treatment ($\pi_{1,1}^{N}$). Dashed lines indicate the linear translation (parallel growths).
@@
All these perspectives reinforce the idea that imposing the parallel growths assumption on the quantities leads to desirable and interpretable restrictions on the evolution of category shares. It links the economic intuition of stable evolution of preferences with a geometric structure of parallel movements in the probability simplex.
@@ 2. Extensions With Multiple Time Periods
In this section, I propose several extensions of the method in the case where we have multiple time periods and possibly more potentially heterogeneous control groups.
2.1. Relaxations of parallel growths
When multiple pre-treatment periods are available, the plausibility of the parallel growths assumption can be evaluated by examining the pre-treatment trajectories of the log-quantities. If the estimated pre-treatment trends appear non-parallel, this raises concerns about the validity of the assumption. Rather than discarding the design altogether, one can relax the assumption in a way that delivers informative bounds on the treatment effect. In particular, I adopt a relaxation strategy closely related to the approach in Ban and Kédagni (2022). They relax the parallel trends assumption by allowing the counterfactual difference between treated and control units to lie within the convex hull of the differences observed in the two most recent pre-treatment periods ($t=0$ and t=1t=-1). By taking the convex hull of the last two differences, they are saying: the counterfactual evolution cannot jump outside the range spanned by recent dynamics. If the differences between groups have been changing slowly, then the most recent differences provide a plausible envelope for what would have happened in the absence of treatment. This is weaker than requiring parallelism and rules out wild deviations inconsistent with observed history. Analogously, in our setting, I extend this idea by applying the same type of convex-hull relaxation. I consider a more general case, where one can either only use the two most recent periods, all the past periods, or all the past periods with higher weights to the most recent ones. I keep the same notation as before, but now, we have that $t\in\{-T_{1},...,0,1\}$.
Assumption 3.
For every category $c_{k}$ where $k=1,\dots,p$, we have:
$\log\left(q_{1,1}^{N}(c_{k})\right)-\log\left({q_{0,1}^{N}(c_{k})}\right)\Bigg[\min_{t\leq 0}\quad\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right),\quad\max_{t\leq 0}\quad\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)\Bigg].$
[…]
@@ 2.1. Relaxations of parallel growths
$GTT\in\left[\frac{\sum_{k=1}^{p}q_{1,1}^{I}(c_{k})}{\sum_{k=1}^{p}b_{max}(c_{k})}-1,\frac{\sum_{k=1}^{p}q_{1,1}^{I}(c_{k})}{\sum_{k=1}^{p}b_{min}(c_{k})}-1\right],$
$CTT\in\left\{\pi\in\mathcal{S}^{p-1}:\pi=\pi^{I}_{1,1}\ominus s,s\in\ell^{-1}\left(\left\{r\in\mathbb{R}^{p-1}\,\Big|\,\log(b_{min}(c_{k}))\leq\ r_{k}\leq\log(b_{max}(c_{k})),\right\}\right)\right\}$
where we have:
$b_{min}(c_{k})=\exp\left(\min_{t\leq 0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)+\log(q_{0,1}^{N}(c_{k}))\right),$
$b_{max}(c_{k})=\exp\left(\max_{t\leq 0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)+\log(q_{0,1}^{N}(c_{k}))\right).$
@@
Theorem 2 provides a clear characterization of how partial identification operates under the relaxed parallel growths assumption. It establishes sharp bounds on both the growth treatment effect on the treated and the compositional treatment effect on the treated by allowing for deviations from parallelism between treated and control groups. The CTT bounds describe a feasible set within the simplex of compositions, capturing all possible redistributions consistent with those counterfactuals. The choice of $\omega_{t}$ governs the degree of relaxation: uniform weights yield a fully agnostic specification based solely on pre-treatment extremes, while time-increasing weights change the bounds by prioritizing near-treatment dynamics.
2.2. Staggered adoption
In this section, I extend the staggered treatment adoption framework of Callaway and Sant’Anna (2021), originally developed for scalar outcomes, to the setting of categorical outcomes. Such designs are increasingly common in applied policy evaluation, where reforms or programs are rolled out gradually across units rather than implemented simultaneously. I consider a panel of $n$ units observed over periods $t=0,1,\dots,T$. Let $D_{i,t}\in\{0,1\}$ indicate whether unit $i$ is treated in period $t$. I impose the standard irreversibility assumption:
Assumption 4 (Irreversibility of Treatment).
No unit is treated at baseline: $D_{i,0}=0$ almost surely. Moreover, treatment is permanent: if $D_{i,t-1}=1$, then $D_{i,t}=1$ for all $t=1,\dots,T$.
Under this assumption, each unit that ever receives treatment has a well-defined first treatment time, denoted $G_{i}=\min\{t:D_{i,t}=1\}$. Units never treated are assigned $G_{i}=\infty$. This partitions the sample into cohorts: for each $g\in\{1,\dots,T\}$, group $G_{g}$ consists of units first treated in period $g$; the never-treated group is $G_{\infty}$. The outcome of interest is categorical: for each group and time, I observe a vector of positive quantities $q_{G_{g},t}$. I also work with the associated share vector $\pi_{G_{g},t}=q_{G_{g},t}/\sum_{k=1}^{p}q_{G_{g},t}(c_{k})$.
For each cohort $G_{g}$, I observe pre-treatment (untreated) outcomes for all $t<g$, and post-treatment (treated) outcomes for $t\geq g$. The goal is to identify the counterfactual untreated quantity $q_{G_{g},t}^{N}$ for $t\geq g$. Let $\bar{g}=\max\{G_{i}:G_{i}<\infty\}$ denote the last period in which any unit is first treated, and define $\mathcal{G}=\text{Supp}(G)\setminus\{\bar{g}\}$ as the set of all treated cohorts except the last one. (The last cohort cannot be used as a control for itself under the second identification strategy, as explained below). I propose two identification strategies, both generalizing the parallel growths assumption (assumption 2) to the staggered setting:
Assumption 5 (Parallel Growths : Never-Treated Control).
For every $g\in\mathcal{G}$ and every $t\geq g$,
$\log(q_{G_{g},t}^{N})-\log(q_{G_{\infty},t}^{N})=\log(q_{G_{g},t-1}^{N})-\log(q_{G_{\infty},t-1}^{N}).$
Assumption 6 (Parallel Growths : Not-Yet-Treated Controls).
For every $g\in\mathcal{G}$, and for all $s,t$ such that $t\geq g$ and $t\leq s<\bar{g}$,
$\log(q_{G_{g},t}^{N})-\log(q_{G_{s},t}^{N})=\log(q_{G_{g},t-1}^{N})-\log(q_{G_{s},t-1}^{N}).$
Assumption 5 states that, in the absence of treatment, the log-quantity paths of cohort $G_{g}$ and the never-treated group $G_{\infty}$ evolve in parallel. Assumption 6 replaces the never-treated group with not-yet-treated cohorts $G_{s}$ (those first treated after time $t$), which serve as internal controls. These conditions enable us to identify counterfactuals by extrapolating pre-treatment log-gaps into post-treatment periods. The resulting identification formulas are constructive, interpretable, and directly estimable from observed data, as formalized in the following theorem.
Theorem 3 (Identification).
Under Assumption 5, for every $g\in\{1,\dots,T\}$ and every $t\geq g$,
$q_{G_{g},t}^{N}=\exp\Big(\log q_{G_{g},g-1}^{N}+\log q_{G_{\infty},t}^{N}-\log q_{G_{\infty},g-1}^{N}\Big).$
Under Assumption 6, for every $g\in\mathcal{G}$ and every $t\geq g$, and for any $s$ such that $t<s\leq\bar{g}$ (i.e., $G_{s}$ is not treated by time $t$),
$q_{G_{g},t}^{N}=\exp\Big(\log q_{G_{g},g-1}^{N}+\log q_{G_{s},t}^{N}-\log q_{G_{s},g-1}^{N}\Big).$
The corresponding distributions are obtained by normalizing the counterfactual quantities and the total by summing them.
2.3. Synthetic CoDiD
Now suppose we have a setting with $G>2$ groups observed over $T>2$ time periods. Only the first group ($g=1$) receives treatment, beginning after period $T_{0}$; all other groups ($g=2,\dots,G$) serve as never-treated controls. In this context, Arkhangelsky et al. (2021) propose the Synthetic difference-in-differences (SDID) method, an extension of the classic DiD framework that constructs a data-driven synthetic control by optimally reweighting the control units to best reproduce the pre-treatment trajectory of the treated unit. This improves the credibility of post-treatment comparisons by enforcing a stronger pre-treatment fit than standard DiD. In this section, I adapt SDID to settings where the outcome is a categorical quantity vector—for example, expenditures across categories, counts of events by type, or any positive-valued composition. For each group $g$ and time $t$, let $q_{g,t}$ denote the vector of quantities across $p$ categories, and let $T_{0}$ be the last pre-treatment period. The objective is to estimate treatment effects on both GTT and CTT. The procedure is applied category by category; for clarity of exposition, I suppress the category index $(c_{k})$ in what follows, noting that all steps are repeated for each component of the quantity vector.
  • Find Unit Weights($\boldsymbol{\omega}^{\ast}$): I find weights for the control groups so that their weighted pre-treatment path best matches the pre-treatment path of the treated group.
[…]
  • Find Time Weights ($\boldsymbol{\lambda}^{\ast}$): I find weights for the pre-treatment time periods so that the pre-treatment average of the controls best matches their post-treatment average.
[…]
The average categorical quantities are obtained by:
$\text{(Treated, Before)}\quad\boldsymbol{q}_{\text{treated,pre}}=\exp\left(\sum_{t=1}^{T_{0}}\lambda^{\ast}_{t}\log(\boldsymbol{q}_{1,t})\right).$
$\text{(Treated, After)}\quad\boldsymbol{q}_{\text{treated,post}}=\exp\left(\sum_{t=T_{0}+1}^{T}\frac{1}{T-T_{0}}\ell(\boldsymbol{q}_{1,t})\right).$
$\text{(Control, Before)}\quad\boldsymbol{q}_{\text{control,pre}}=\exp\left(\sum_{g=2}^{G}\sum_{t=1}^{T_{0}}\omega^{\ast}_{j}\lambda^{\ast}_{t}\log(\boldsymbol{q}_{g,t})\right).$
$\text{(Control, After)}\quad\boldsymbol{q}_{\text{control,post}}=\exp\left(\sum_{g=2}^{G}\sum_{t=T_{0}+1}^{T}\omega^{\ast}_{j}\frac{1}{T-T_{0}}\log(\boldsymbol{q}_{g,t})\right).$
Finally, I calculate the average growth treatment effect.
$\text{GTT}=\frac{\boldsymbol{q}_{\text{treated,post}}}{\boldsymbol{q}_{\text{treated,pre}}+\boldsymbol{q}_{\text{control,post}}-\boldsymbol{q}_{\text{control,pre}}}-1.$
We then construct the distributions by normalizing across categories: $\boldsymbol{\pi}_{\text{treated,pre}}$, $\boldsymbol{\pi}_{\text{treated,post}}$, $\boldsymbol{\pi}_{\text{control,pre}}$, $\boldsymbol{\pi}_{\text{control,post}}$.
The Compositional Treatment Effect (CTT) is defined as:
$\text{CTT}=\left(\boldsymbol{\pi}_{\text{treated,post}}\ominus\boldsymbol{\pi}_{\text{control,post}}\right)\oplus\left(\boldsymbol{\pi}_{\text{treated,pre}}\ominus\boldsymbol{\pi}_{\text{control,pre}}\right).$
This effect shows how the treatment changed the probability distribution for the treated group, relative to the change experienced by the synthetic control group. It tells us which categories gained or lost probability mass because of the treatment.
3. Connexions With Existing Approaches
In this section, I discuss the link between my method and existing methods in the literature.
3.1. Connexion with Logistic model
Recall: $Y_{it}\in\{c_{1},c_{2},\dots,c_{p}\}$: outcome for unit $i$ at time $t$, $G_{i}\in\{0,1\}$: group indicator (1 = treated), $T_{t}\in\{0,1\}$: time indicator (1 = post), Interaction: $D_{it}=G_{i}\cdot T_{t}$ (standard DiD treatment indicator). Pick category $c_{p}$ as the baseline. The two-way fixed effects multinomial logit model without covariates is:
$\mathbb{P}(Y_{it}=c_{k}\mid G_{i},T_{t})=\frac{\exp\left(\alpha_{k}+\delta_{k}G_{i}+\gamma_{k}T_{t}+\beta_{k}D_{it}\right)}{1+\sum_{g=1}^{p-1}\exp\left(\alpha_{j}+\beta_{j}G_{i}+\gamma_{j}T_{t}+\delta_{j}D_{it}\right)},\quad k=1,\dots,p-1.$
This specification is fully saturated: for each non-baseline outcome category $c_{k}$, the model includes four parameters, an intercept $\alpha_{k}$, a group effect $\delta_{k}$, a time effect $\gamma_{k}$, and an interaction term $\beta_{k}$, which together exactly span the log-odds in all four group–time cells. The interaction coefficient $\beta_{k}$ captures the difference-in-differences contrast in log-odds and admits a causal interpretation under the assumption of parallel trends on the log-odds. The compositional treatment effect parameter (CTT) can be linked to the coefficients of a multinomial logit model as follows:
$CTT=\ell^{-1}\big(\beta_{1},\beta_{2},\dots,\beta_{p-1}\big),$
where $\ell^{-1}(\cdot)$ denotes the inverse compositional log-ratio transformation, which maps the vector of log-odds coefficients back into category shares. Although using the multinomial logit representation coincides with my approach to modeling category shares in the simple $(2\times 2)$ setup, several conceptual and practical differences are worth emphasizing.
The first key difference is that the multinomial logit model does not directly yield counterfactual quantities without imposing additional structural assumptions. In contrast, my approach provides these counterfactuals constructively, enabling the recovery of both treatment effects on quantities and on shares within a unified framework. Second, in settings with multiple time periods, the multinomial logit specification offers no straightforward way to relax the identifying assumptions when they appear violated. My framework, however, accommodates such relaxations naturally, formalized through assumptions 3 allowing for partial identification. Third, when treatment timing varies across units (staggered adoption), the multinomial logit model requires estimating a full vector of time-varying treatment effect coefficients. By contrast, my method allows the estimation of treatment effect for each post-treatment period independently, without the need for joint optimization or functional minimization. This feature makes my approach computationally lighter. Finally, the Compositional Treatment Effect (CTT) parameter provides a direct and interpretable measure of the treatment’s impact on category shares. Operating within the space of compositional data, it connects naturally to discrete choice theory and can be naturally used in applications where proportions or probabilities, rather than levels, constitute the primary outcomes of interest.
3.2. Unified view with DiD and CiC
My method identifies the counterfactual probability mass function for unordered categorical outcomes. I show how the identifying assumption relates to parallel trends (DiD) and rank invariance (CiC). DiD targets objects like expectations, CiC targets the CDF for ordered outcomes, and CoDiD extends the same logic to the PMF for unordered outcomes. I proceed in two steps. First, I show that DiD and CiC, although targeting different objects (means versus CDFs), share the same geometric identification logic. Second, I show how CoDiD applies that logic to probability mass functions. A recall of the setup: I consider two groups $g\in{0,1}$ (treatment and control), observed in two periods $t\in{0,1}$. Let $m_{g,t}^{N}$ denote the potential parameter of interest for unit $g$ at time $t$ in the absence of treatment.
In DiD, the parameter of interest $m_{g,t}^{N}$ is the expectation of some individual-level outcome, $m_{g,t}^{N}=E(Y^{N}_{g,t})$. The construction of the counterfactual is given by:
$E(Y^{N}_{1,1})=E(Y^{N}_{1,0})\underbrace{-E(Y^{N}_{0,0})+E(Y^{N}_{0,1}).}_{\text{control group change over time}}$
In CiC, the parameter of interest $m_{g,t}^{N}$ is the cumulative distribution function of some individual-level outcome, $m_{g,t}^{N}=F_{Y^{q}_{g,t}}$. The construction of the counterfactual is given by:
$F_{Y^{N}_{1,1}}=F_{Y^{N}_{1,0}}\circ\underbrace{F^{-1}_{Y^{N}_{0,0}}\circ F_{Y^{N}_{0,1}}.}_{\text{control group change over time}}.$
DiD and CiC share a simple underlying idea: the object of interest exists within a space where the notion of movement can be well-defined, and there is a transformation that characterizes how the control group evolves within this space. To obtain the counterfactual for the treated group, we then apply the same transformation to the treated group. Mathematically, if the control group moves from $m_{0,0}^{N}$ to $m_{0,1}^{N}$, I write:
$m_{0,1}^{N}=\varphi^{c}(m_{0,0}^{N}),$
where $\varphi^{c}$ captures the change or movement in the control group. The counterfactual for the treated group is then obtained by applying the same transformation to the initial state of the treated group $m_{1,0}^{N}$:
$m_{1,1}^{N}=\varphi^{c}(m_{1,0}^{N}).$
In DiD:
$\forall m\in\mathbb{R},\quad\varphi^{c}(m)=m-E_{Y^{N}_{0,0}}+E_{Y^{N}_{0,1}}.$ (12)
In CiC:
$\forall F\in\mathcal{F},\quad\varphi^{c}(F)=F\circ F^{-1}_{Y^{N}_{0,0}}\circ F_{Y^{N}_{0,1}}.$ (13)
where $\mathcal{F}$ is the space of cumulative distribution functions. The map $\varphi^{c}$ has the same interpretation in both methods and can be understood as a translation map. If I want to use a map to define the idea of movement within a space. A natural way to formalize that is through the concept of translation. A translation operator is defined as an operator that shifts points by a certain amount in a certain direction. I consider a general definition of this operator. Intuitively, suppose the outcome space forms a group, meaning it is a set of objects equipped with a binary operation that can help define the idea of movement (moves that can be combined, reversed, and have a neutral element that does nothing). These moves can be addition (for numeric outcomes) or function composition (for distributional transformations). Let’s start with the formal definition of a group.
Definition 1 (Group).
A group $(G,\wedge)$ consists of a set $G$ and a binary operation
$\wedge:G\times G\to G,\quad(a,b)\mapsto a\wedge b,$
satisfying the following axioms: i) Associativity: $(a\wedge b)\wedge c=a\wedge(b\wedge c)$ for all $a,b,c\in G$. ii) Identity: There exists $e\in G$ such that $e\wedge a=a\wedge e=a$ for all $a\in G$. iii) Inverses: For each $a\in G$, there exists $a^{-1}\in G$ such that $a^{-1}\wedge a=a\wedge a^{-1}=e$.
For example, in DiD, the outcome space is $\mathbb{R}$ with addition ++ as the group operation, identity $e=0$, and inverse given by negation a-a. In CiC, the outcome space is the set of cumulative distribution functions, the operation is composition $\circ$, the identity is the identity function $Id(x)=x$, and the inverse is the quantile function. Once the group structure is defined, I can formalize the idea of a movement using translations.
Definition 2 (Right Translation Operator (Hamilton, 2017)).
Let $(G,\wedge)$ be a group with identity element $e$. For a fixed $h\in G$, the left translation by $h$ is the map:
$R_{h}:G\to G,\qquad R_{h}(m):=m\wedge h,\quad\forall h\in G.$
If you start at $m$, you then apply $h$ on the right, ending at $m\wedge h$. Think of shifting relative to where you already are, like taking the current position and then performing the move locally. This formalizes the notion of a shift or transformation acting consistently across the entire outcome space. From this perspective, the map $\varphi^{c}$ can be identified as the unique right translation that captures how the control group evolves over time.
In DiD, $\varphi^{c}$ is the right translation operator defining how the control group has moved from its initial state. we have that:
$\varphi^{c}(m)=R_{h}\left(m\right)=m+h.$
Since $\varphi^{c}(E(Y^{N}_{0,0}))=E(Y^{N}_{0,1})$, it implies that $h=E(Y^{N}_{0,1})-E(Y^{N}_{0,0}).$ We therefore obtain that:
$\varphi^{c}(m)=m+E(Y^{N}_{0,1})-E(Y^{N}_{0,0}).$
Graphically, it can be represented in figure 4
[…]
Figure 4. Geometric Illustration of parallel trends
In CiC, $\varphi^{c}$ is also the left translation operator defining how the control group has moved from its initial state. we have that:
$\varphi^{c}(F)=R_{h}\left(F\right)=F\circ h.$
Since $\varphi^{c}(F_{Y^{N}_{0,0}})=F_{Y^{N}_{0,1}}$, it implies that $h=F^{-1}_{Y^{N}_{0,0}}\circ F_{Y^{N}_{0,1}}$. We therefore obtain that:
$\varphi^{c}(F)=F\circ F^{-1}_{Y^{N}_{0,0}}\circ F_{Y^{N}_{0,1}}.$
Graphically, I can illustrate through figure 5.
$\mathcal{F}$$F_{Y^{N}_{0,0}}$$F_{Y^{N}_{0,1}}$$F_{Y^{N}_{1,0}}$$F_{Y^{N}_{1,1}}$
Figure 5. Geometric illustration of Rank invariance
The uniqueness of $\varphi^{c}$ is what guarantees point identification. If multiple transformations could map the control group across periods, the counterfactual is no longer unique. In that case, we have identified a set of feasible outcomes, as in the case of Changes-in-Changes (CiC) with a discrete ordered outcome.
CoDiD shares this same logic, but for probability mass functions, where the space is the simplex. Here, the parameter of interest $m_{g,t}^{N}$ is the probability mass function, $m_{g,t}^{N}=\pi^{N}_{g,t}$. The construction of the counterfactual is given by:
$\pi^{N}_{1,1}=\pi^{N}_{1,0}\underbrace{\ominus\pi^{N}_{0,0}\oplus\pi^{N}_{0,1}.}_{\text{control group change over time}}$
The control group map is therefore defined in this way:
$\forall\pi\in\mathcal{S}^{p-1},\quad\varphi^{c}(\pi)=\pi\ominus\pi^{N}_{0,0}\oplus\pi^{N}_{0,1}.$ (14)
In CoDiD, the group is the simplex equipped with the perturbation operator $(\oplus)$. In that method, $\varphi^{c}$ is also a right translator operator defined as:
$\varphi^{c}(\pi)=R_{h}\left(\pi\right)=\pi\oplus h.$
Since $\varphi^{c}(\pi^{N}_{0,0})=\pi^{N}_{0,1}$, it implies that $h=\ominus\pi^{N}_{0,0}\oplus\pi^{N}_{0,1}$. We therefore obtain that:
$\varphi^{c}(\pi)=\pi\ominus\pi^{N}_{0,0}\oplus\pi^{N}_{0,1}$
That unified logic breaks down if I use the Euclidean group structure for the simplex (doing parallel trends on proportions). In fact, when I restrict to the simplex, the structure no longer forms a group since the neutral element would have to be the zero vector (which is not part of the simplex). In contrast, under my formulation, the group structure is preserved, ensuring consistency with DiD and CiC. Table 1 summarizes this comparison. Let $\mathcal{F}$, be the space of non-decreasing real functions.
Table 1. Comparison of Identification Strategies
Method Outcome Type Geometric Space Control group map
parallel trends (DiD) Expectation $\mathbb{R}$ (Euclidean) $m\mapsto m-E_{Y^{N}_{0,0}}+E_{Y^{N}_{0,1}}$
Rank Invariance (CiC) CDF $\mathcal{F}$ $F\mapsto F\circ F^{-1}_{Y^{N}_{0,0}}\circ F_{Y^{N}_{0,1}}$
parallel growths (CoDiD) PMF Aitchison simplex $\pi\mapsto\pi\ominus\pi^{N}_{0,0}\oplus\pi^{N}_{0,1}$
4. Applications
This section presents two empirical applications. The first examines the effect of early voting policies on voter support decisions in U.S. presidential elections. The second evaluates the impact of the regional greenhouse gas initiative (RGGI) on the composition of electricity generation across energy sources.
4.1. Impact of early voting law on distributions of voters’ support across parties

### v3

@@ 1.2.3. Geometric justifications of parallel growths
In this section, I provide some additional geometric justifications for parallel growths. I start by interpreting the restriction on the shares using the compositional difference operator.
@@
Geometric interpretation: Parallel growths also imply parallel trajectories of shares in the probability simplex. shares lie in the simplex $\mathcal{S}^{p-1}$, where movement can be meaningfully defined using the geometry introduced by Aitchison (1982). Unlike the Euclidean case, this geometry accounts for the simplex’s constraints and forms the basis of compositional data analysis, redefining operations such as addition, scaling, and distance.
@@
. Let $\mathcal{S}^{p-1}$ be the interior of the p1p-1 simplex, define the following operations: for all $\pi_{1},\pi_{2}\in\mathcal{S}^{p-1}$, we have:
@@
Figure 2. Illustration of parallel growths in the simplex. The red curve represents the trajectory of the control group from pre-treatment ($\pi_{0,0}^{0}$) to post-treatment ($\pi_{0,1}^{0}$), while the blue curve shows the counterfactual trajectory of the treated group from pre-treatment ($\pi_{1,0}^{0}$) to post-treatment ($\pi_{1,1}^{0}$). Dashed lines indicate the linear translation (parallel growths).
@@
To sum up, all these perspectives reinforce the idea that imposing the parallel growths assumption on the quantities leads to desirable and interpretable restrictions on the evolution of category shares and totals. It links the economic intuition of stable evolution of preferences with a geometric structure of parallel movements in the probability simplex.
@@ 2. Extensions With Multiple Time Periods
When multiple pre-treatment periods are available, the plausibility of the parallel growths assumption can be evaluated by examining the pre-treatment trajectories of the log-quantities. If the estimated pre-treatment trends appear non-parallel, this raises concerns about the validity of the assumption. Rather than discarding the design altogether, one can relax the assumption in a way that delivers informative bounds on the treatment effect. In particular, I adopt a relaxation strategy closely related to the approach in Ban and Kédagni (2022). They relax the parallel trends assumption by allowing the counterfactual difference between treated and control units to lie within the convex hull of the differences observed in the two most recent pre-treatment periods ($t=0$ and t=1t=-1). By taking the convex hull of the last two differences, they are saying: the counterfactual evolution cannot jump outside the range spanned by recent dynamics. If the differences between groups have been changing slowly, then the most recent differences provide a plausible envelope for what would have happened in the absence of treatment. This is weaker than requiring parallelism and rules out wild deviations inconsistent with observed history. Analogously, in our setting, I extend this idea by applying the same type of convex-hull relaxation. I consider a more general case, where one can either only use the two most recent periods, all the past periods, or all the past periods with higher weights to the most recent ones. I keep the same notation as before, but now, we have that $t\in\{-T_{1},...,0,1\}$. First, define the pre-treatment log-differences between groups for each category $k=1,\dots,p$ and take their minimum and maximum.
$d^{\min}_{k}=\min_{t\leq 0}\Bigg(\log\left(q^{0}_{k,1,t}\right)-\log\left(q^{0}_{k,0,t}\right)\Bigg),\quad d^{\max}_{k}=\max_{t\leq 0}\Bigg(\log\left(q^{0}_{k,1,t}\right)-\log\left(q^{0}_{k,0,t}\right)\Bigg)$
Assumption 3 (Robust Parallel Growth).
For each category $k$, the post-treatment change in log-quantities between treated and control groups is bounded by the range of pre-treatment differences:
$d_{k,1}=\log\left(q^{0}_{k,1,1}\right)-\log\left(q^{0}_{k,0,1}\right)\in\big[d^{\min}_{k},d^{\max}_{k}\big].$
This assumption relaxes the strict parallel growth requirement by allowing the post-treatment log-change to vary within the historical range observed in pre-treatment periods. It ensures that the counterfactual outcome for each category remains plausible, respecting the variability seen in the pre-treatment trends. When $d^{\min}_{k}=d^{\max}_{k}$, the pre-treatment trends are exactly parallel, and the standard parallel growth assumption is recovered. Practically, the assumption can be implemented in three steps: compute pre-treatment differences, find their minimum and maximum, and constrain the post-treatment log-change to lie within this interval. This approach yields bounds on counterfactual quantities rather than point estimates, providing a transparent measure of uncertainty when strict parallel growth is questionable. It preserves the interpretability of the DiD framework, as the treated group’s counterfactual is still anchored to the evolution of the control group, but now allows for more flexible deviations observed historically. It is also possible to add weights for each time period to determine how much each pre-treatment period contributes to this range. In practice, more recent pre-treatment periods can receive greater importance. The identified set is the collection of all quantities that are consistent with the data and the model assumptions.
@@
$b_{k}^{\min}=\exp\!\left(d^{\min}_{k}+\log q^{0}_{k,0,1}\right),\quad b_{k}^{\max}=\exp\!\left(d^{\max}_{k}+\log q^{0}_{k,0,1}\right)$
$q^{0}_{k,1,1}\in[\,b_{k}^{\min},\;b_{k}^{\max}\,]$
$S^{0}_{1,1}\in\Bigg[\;\sum_{k=1}^{p}b_{k}^{\min},\;\sum_{k=1}^{p}b_{k}^{\max}\;\Bigg]$
$\pi^{0}_{1,1}\in\left\{\pi\in\mathcal{S}^{p-1}:\pi_{k}=\frac{q_{k}}{\sum_{j}q_{j}},\;q_{k}\in[b_{k}^{\min},b_{k}^{\max}]\right\}.$
@@
Theorem 2 establishes sharp bounds on both the growth treatment effect on the treated and the compositional treatment effect on the treated by allowing for deviations from parallelism between treated and control groups. The CTT bounds describe a feasible set within the simplex of compositions, capturing all possible redistributions consistent with those counterfactuals. The choice of $\omega_{t}$ governs the degree of relaxation: uniform weights yield a fully agnostic specification based solely on pre-treatment extremes, while time-increasing weights change the bounds by prioritizing near-treatment dynamics.
3. Application: Impact of early voting law on distributions of voters’ support across parties